An electrolytic capacitor is charged to 15 V when a constant current of 1.0 A is passed through it for 30 µs. What is the value of its capacitance?     


1
Expert's answer
2022-02-28T13:44:16-0500

"Charge, Q=It=1\times 30\times 10^{-6}=3\times 10^{-5}C"

Capacitance, "C=Q/V=\frac{3\times 10^{-5}}{15}=2.0 \mu F"